# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 4.4 Global stability
The perturbed solution exists in all of $[s_{0},+\infty)\times\Sigma_{s}$, satisfying the global estimate:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{2Hs},\qquad\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs},\qquad\mathcal{E}_{N}(s)\leq C\mathring{\varepsilon}^{2},$ (4.19)
@@
$\mathcal{E}_{N}(s)\leq C\bigg{[}\mathcal{E}_{N}(s_{0})+\int^{s_{b}}_{s_{0}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big{\}}\mathrm{d}s\bigg{]}$
@@
$\mathcal{E}_{N}(s)\leq$ $\,C\bigg{[}\mathring{\varepsilon}^{2}+\int^{s_{*}}_{s_{0}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}ds$
$+\int^{s_{b}}_{s_{*}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big{\}}ds\bigg{]}$
$\leq$ $\,C\big{[}\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}\big{]}$
@@ 5 Future stability estimates
Our goal in this section is to derive the energy estimates that prove Theorem 2. We derive the higher order equations in Section 5.2, and treat the error estimates in Section 5.3. The overall energy estimate is proven in Proposition 5.10 at the end of Section 5.4. We start as an introduction with the basic energy identity in this gauge.
@@ 5.1 Discussion of the energy identity
In the currrent setting, the first variation equation (2.21) can also be written as:
@@
Similarly for $g^{-1}$. Indeed, it follows from (2.25) that
$\partial_{s}g^{ij}+2Hg^{ij}=$ $\,(\mathrm{Error}_{g^{-1}})^{ij}$ (5.5)
$(\mathrm{Error}_{g^{-1}})^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\Phi g^{ia}\bigl{(}\widetilde{k}_{a}{}^{j}-H\delta_{a}{}^{j}\Phi^{-1}\bigr{)}$ (5.6)
@@
$\frac{1}{2}\partial_{s}\bigl{(}e^{4Hs}\widehat{\Phi}^{2}\bigr{)}+e^{4Hs}|\nabla\widehat{\Phi}|^{2}_{g}=e^{4Hs}\nabla^{i}\bigl{(}\widehat{\Phi}\partial_{i}\widehat{\Phi}\bigr{)}+(\mathrm{Error}_{\widehat{\Phi}})e^{4Hs}\widehat{\Phi}$ (5.9)
@@
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g_{cj}\Phi\nabla^{a}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma}})_{ic}^{a}$ (5.10a)
@@
$+(\mathrm{Error}_{\widehat{k}})_{i}{}^{j}$
@@
$\frac{1}{2}\partial_{s}\Bigl{(}e^{2Hs}|\widehat{k}|_{g}^{2}+\frac{1}{3}e^{2Hs}|\widehat{\Gamma}|_{g}^{2}\Bigr{)}+2H|\widehat{k}|_{g}^{2}=\frac{2}{3}\Phi g^{ii^{\prime}}g_{jj^{\prime}}\nabla^{a}\bigl{(}e^{2Hs}\widehat{\Gamma}_{ai^{\prime}}^{j^{\prime}}\widehat{k}_{i}{}^{j}\bigr{)}-\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{a}\bigl{(}e^{2Hs}\widehat{\Gamma}^{a}_{ij}\widehat{k}_{i^{\prime}}{}^{j}\bigr{)}\\ +e^{2Hs}\widehat{k}_{i}{}^{j}\nabla^{i}\nabla_{j}\widehat{\Phi}+\frac{1}{3}\Phi e^{2Hs}g^{ii^{\prime}}\widehat{k}_{i}{}^{j}\nabla_{j}\widehat{\Gamma}_{i^{\prime}a}^{a}-\frac{2}{3}\Phi e^{2Hs}g^{ab}g_{jj^{\prime}}\widehat{k}_{i}{}^{j}\nabla^{i}\widehat{\Gamma}_{ab}^{j^{\prime}}\\ +e^{2Hs}\mathrm{Error}_{g,g^{-1}}\star(\widehat{\Gamma}\star\widehat{\Gamma}+\widehat{k}\star\widehat{k})+e^{2Hs}\mathrm{Error}_{\widehat{\Gamma}}\star\widehat{\Gamma}$ (5.14)
@@
$\frac{1}{2}\partial_{s}\Bigl{(}e^{2Hs}|\widehat{k}|_{g}^{2}+\frac{1}{3}e^{2Hs}|\widehat{\Gamma}|_{g}^{2}\Bigr{)}+2H|\widehat{k}|_{g}^{2}=\Phi\>\mathrm{div}_{g}(e^{2Hs}\widehat{\Gamma}\star\widehat{k})+\nabla^{i}\Bigl{(}e^{2Hs}\widehat{k}_{i}{}^{j}\partial_{j}\widehat{\Phi}\Bigr{)}\\ -e^{2Hs}|\nabla\widehat{\Phi}|_{g}^{2}+\frac{1}{3}\Phi e^{2Hs}\nabla^{i}\widehat{\Phi}\>\widehat{\Gamma}_{ia}^{a}-\frac{2}{3}\Phi e^{2Hs}g^{ab}\nabla_{j}\widehat{\Phi}\>\widehat{\Gamma}_{ab}^{j}\\ +e^{2Hs}\mathrm{Error}_{\mathrm{div}\widehat{k}}\star\nabla\widehat{\Phi}+e^{2Hs}\mathrm{Error}_{\mathrm{div}\widehat{k},\widehat{\Gamma}}\star\widehat{\Gamma}+e^{2Hs}\mathrm{Error}_{g,g^{-1}}\star(\widehat{\Gamma}\star\widehat{\Gamma}+\widehat{k}\star\widehat{k})$ (5.17)
@@
$\mathcal{E}(s)=\,\|\widehat{g}\|^{2}_{L^{2}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{L^{2}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{L^{2}(\Sigma_{s},g)}+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{L^{2}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{L^{2}(\Sigma_{s},g)}\big{)}$ (5.18)
@@ 5.2.1 Higher order equations
$(\mathrm{Error}_{\widehat{g},\ell})_{ij}=$ ()$\,\nabla^{(\ell)}\big{\{}2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}\big{\}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}_{ij},$ (5.20a)
$(\mathrm{Error}_{\widehat{g}^{-1},\ell})^{ij}=$ ()$-\nabla^{(\ell)}\big{\{}2\Phi g^{ia}\widehat{k}_{a}{}^{j}+2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}+2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\big{\}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}^{ij}\,.$ (5.20b)
@@
These equations result by commuting the equations (2.37), (2.38), (2.45) with $\nabla^{(\ell)}$. ∎
@@
$\begin{split}\partial_{s}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}+3H\nabla^{(\ell)}\widehat{k}_{i}{}^{j}=&\,\frac{1}{3}\Phi g^{cj}(\nabla_{c}\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}-\nabla_{a}\nabla^{(\ell)}\widehat{\Gamma}_{ci}^{a})+(\mathrm{Error}_{\widehat{k},\ell})_{i}{}^{j}\\ &+\frac{2}{3}\Phi g^{ab}(\nabla_{a}\nabla^{(\ell)}\widehat{\Gamma}^{j}_{bi}-\nabla_{i}\nabla^{(\ell)}\widehat{\Gamma}^{j}_{ab})+g^{cj}\nabla_{i}\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\end{split}$ (5.23)
@@
$(\mathrm{Error}_{\widehat{k},\ell})_{i}{}^{j}=$ ()$\,\nabla^{(\ell)}\big{\{}\mathfrak{K}_{i}{}^{j}-\widetilde{\Phi}(\widetilde{k}_{l}{}^{l}-3H\widetilde{\Phi}^{-1})\widehat{k}_{i}{}^{j}+(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\big{\}}$ (5.24)
@@
$\partial_{s}\nabla^{(\ell)}\widehat{\Gamma}_{ic}^{a}=\Phi\nabla_{i}\nabla^{(\ell)}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\nabla^{(\ell)}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma},\ell})_{ic}^{a}$ (5.25)
@@
$-g^{ab}g_{cj}\nabla^{(\ell_{1})}\Phi\nabla^{(\ell_{2})}\nabla_{b}\widehat{k}_{i}{}^{j}\big{\}}$
@@
$(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})_{i}=$ (){~$\,\nabla^{(\ell)}\big{\{}\widetilde{\mathfrak{C}}_{i}-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})\big{\}}$ (5.28a)
@@
$(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{j}=$ ()$\,\nabla^{(\ell)}\big{\{}\widehat{g}^{jc}\widetilde{\nabla}_{c}(\widetilde{k}_{l}{}^{l}-3H)-\widehat{g}^{im}\widetilde{\nabla}_{m}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})-g^{im}\widehat{\Gamma}_{mc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})$ (5.28b)
@@ 5.2.2 Higher order energy identities
we use (2.25) and (5.19a) and obtain
$\begin{split}\frac{1}{2}\partial_{s}\big{\{}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\big{\}}=&\,\ell He^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\\ &-\Phi k_{a}{}^{b_{1}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}a}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ &-\ldots-\Phi k_{a}{}^{b_{\ell}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}a}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{i^{\prime}}e^{2\ell Hs}g^{ia}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{j^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{ja}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ +e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}&g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\big{\{}2H\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}+(\mathrm{Error})_{\widehat{g},\ell})_{ij}\big{\}}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ \end{split}$ (5.31)
@@
$\Phi k_{a}^{b}=\Phi\widehat{k}_{a}{}^{b}+\Phi\widetilde{k}_{a}{}^{b}-H\delta_{a}{}^{b}+H\delta_{a}{}^{b}$ (5.32)
@@
$\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell}=e^{2Hs}e^{2\ell Hs}\bigg{[}\frac{2}{3}\Phi g^{ii^{\prime}}g^{cc^{\prime}}g_{aa^{\prime}}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}\star\nabla^{(\ell)}\widehat{k}_{c}{}^{a}\big{]}\\ -\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{b}\big{[}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}j}^{b}\star\nabla^{(\ell)}\widehat{k}_{i}{}^{j}\big{]}+\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{c}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big{]}\\ -\frac{2}{3}\Phi g^{ab}g^{ii^{\prime}}g_{jj^{\prime}}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ab}^{j}\big{]}+g^{ii^{\prime}}g_{jj^{\prime}}g^{cj}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\big{]}\bigg{]}$ (5.36)
where the $\star$ symbol in all schematic expressions above signifies that all relevant indices in these terms are contracted.
@@
(5.25)×$\times\frac{1}{3}e^{2Hs}e^{2\ell Hs}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}$
@@
It remains to differentiate by parts the terms in the RHS which contain $\ell+1$ spatial derivatives of $\widehat{\Gamma}$ and $\widehat{k}$. This produces on one hand the divergence terms $(\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell})$ in (5.36), and on the other hand, divergences of $\nabla^{(\ell)}k$ which we can replace using the higher order constraint equations (5.27). For instance,

### v2

@@ 4.4 Global stability
The perturbed solution exists in all of $s_{0},+\infty)\times\Sigma_{s}$, satisfying the global estimate:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{2Hs},\hskip 18.49988pt\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs},\hskip 18.49988pt\mathcal{E}_{N}(s)\leq C\mathring{\varepsilon}^{2},$ (4.19)
@@
$\mathcal{E}_{N}(s)\leq C\bigg[\mathcal{E}_{N}(s_{0})+\int^{s_{b}}_{s_{0}}e^{\frac{7}{2}Hs}\big\{\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big\}\mathrm{d}s\bigg]\,.$
@@
$\mathcal{E}_{N}(s)\leq$ $\,C\bigg[\mathring{\varepsilon}^{2}+\int^{s_{*}}_{s_{0}}e^{\frac{7}{2}Hs}\big\{\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\mathrm{d}s$
$+\int^{s_{b}}_{s_{*}}e^{\frac{7}{2}Hs}\big\{\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big\}\mathrm{d}s\bigg]$
$\leq$ $\,C\big[\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}\big]\,.$
@@
Remark 4.11 (Local well-posedness).
In the proof of the previous corollary, we tacitly used the local well-posedness of the Einstein vacuum equations in the parabolic gauge (2.31), which is proven independently in Appendix A.
@@ 5 Future stability estimates
Our goal in this section is to derive the energy estimates that prove Theorem 4.1. We derive the higher order equations in Section 5.2, and treat the error estimates in Section 5.3. The overall energy estimate is proven in Proposition 5.10 at the end of Section 5.4. We start as an introduction with the basic energy identity in this gauge.
@@ 5.1 Discussion of the energy identity
In the currrent setting, the first variation equation (2.22a) can also be written as:
@@
Similarly for $g^{-1}$. Indeed, it follows from (2.24b) that
$\partial_{s}g^{ij}+2Hg^{ij}=$ $\,(\mathrm{Error}_{g^{-1}})^{ij}\,$, (5.5)
$(\mathrm{Error}_{g^{-1}})^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\Phi g^{ia}\bigl(\widetilde{k}_{a}{}^{j}-H\delta_{a}{}^{j}\Phi^{-1}\bigr)\,.$ (5.6)
@@
$\frac{1}{2}\partial_{s}\bigl(e^{4Hs}\widehat{\Phi}^{2}\bigr)+e^{4Hs}|\nabla\widehat{\Phi}|^{2}_{g}=e^{4Hs}\nabla^{i}\bigl(\widehat{\Phi}\partial_{i}\widehat{\Phi}\bigr)+(\mathrm{Error}_{\widehat{\Phi}})e^{4Hs}\widehat{\Phi}\,$. (5.9)
@@
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g_{cj}\Phi\nabla^{a}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma}})_{ic}^{a}\,$, (5.10a)
@@
$+(\mathrm{Error}_{\widehat{k}})_{i}{}^{j}\,$.
@@
$\frac{1}{2}\partial_{s}\Bigl(e^{2Hs}|\widehat{k}|_{g}^{2}+\frac{1}{3}e^{2Hs}|\widehat{\Gamma}|_{g}^{2}\Bigr)+2H|\widehat{k}|_{g}^{2}=\frac{2}{3}\Phi g^{ii^{\prime}}g_{jj^{\prime}}\nabla^{a}\bigl(e^{2Hs}\widehat{\Gamma}_{ai^{\prime}}^{j^{\prime}}\widehat{k}_{i}{}^{j}\bigr)-\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{a}\bigl(e^{2Hs}\widehat{\Gamma}^{a}_{ij}\widehat{k}_{i^{\prime}}{}^{j}\bigr)\\ +e^{2Hs}\widehat{k}_{i}{}^{j}\nabla^{i}\nabla_{j}\widehat{\Phi}+\frac{1}{3}\Phi e^{2Hs}g^{ii^{\prime}}\widehat{k}_{i}{}^{j}\nabla_{j}\widehat{\Gamma}_{i^{\prime}a}^{a}-\frac{2}{3}\Phi e^{2Hs}g^{ab}g_{jj^{\prime}}\widehat{k}_{i}{}^{j}\nabla^{i}\widehat{\Gamma}_{ab}^{j^{\prime}}\\ +e^{2Hs}\mathrm{Error}_{g,g^{-1}}\star(\widehat{\Gamma}\star\widehat{\Gamma}+\widehat{k}\star\widehat{k})+e^{2Hs}\mathrm{Error}_{\widehat{\Gamma}}\star\widehat{\Gamma}\,$. (5.14)
@@
[…]
@@
[…]
@@ 5.2.1 Higher order equations
$(\mathrm{Error}_{\widehat{g},\ell})_{ij}=$ $\,\nabla^{(\ell)}\big\{2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}\big\}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}_{ij},$ (5.20a)
$(\mathrm{Error}_{\widehat{g}^{-1},\ell})^{ij}=$ $-\nabla^{(\ell)}\big\{2\Phi g^{ia}\widehat{k}_{a}{}^{j}+2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}+2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\big\}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}^{ij}\,.$ (5.20b)
@@
These equations result by commuting the equations (2.35a), (2.35b), (2.42) with $\nabla^{(\ell)}$. ∎
@@
[…]
@@
$(\mathrm{Error}_{\widehat{k},\ell})_{i}{}^{j}$= $\,\nabla^{(\ell)}\big\{\mathfrak{K}_{i}{}^{j}-\widetilde{\Phi}(\widetilde{k}_{l}{}^{l}-3H\widetilde{\Phi}^{-1})\widehat{k}_{i}{}^{j}+(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\big\}$ (5.24)
@@
$\partial_{s}\nabla^{(\ell)}\widehat{\Gamma}_{ic}^{a}=\Phi\nabla_{i}\nabla^{(\ell)}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\nabla^{(\ell)}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma},\ell})_{ic}^{a}\,$, (5.25)
@@
$-g^{ab}g_{cj}\nabla^{(\ell_{1})}\Phi\nabla^{(\ell_{2})}\nabla_{b}\widehat{k}_{i}{}^{j}\big\}\,.$
@@
[…]
@@
[…]
@@ 5.2.2 Higher order energy identities
we use (2.24b) and (5.19a) and obtain
$\begin{split}\frac{1}{2}\partial_{s}\big\{e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\big\}=&\,\ell He^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\\ &-\Phi k_{a}{}^{b_{1}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}a}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ &-\ldots-\Phi k_{a}{}^{b_{\ell}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}a}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{i^{\prime}}e^{2\ell Hs}g^{ia}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{j^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{ja}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ +e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}&g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\big\{2H\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}+(\mathrm{Error})_{\widehat{g},\ell})_{ij}\big\}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\,.\end{split}$ (5.31)
@@
$\Phi k_{a}^{b}=\Phi\widehat{k}_{a}{}^{b}+\Phi\widetilde{k}_{a}{}^{b}-H\delta_{a}{}^{b}+H\delta_{a}{}^{b}\,$, (5.32)
@@
$\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell}=e^{2Hs}e^{2\ell Hs}\bigg[\frac{2}{3}\Phi g^{ii^{\prime}}g^{cc^{\prime}}g_{aa^{\prime}}\nabla_{i}\big[\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}\star\nabla^{(\ell)}\widehat{k}_{c}{}^{a}\big]\\ -\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{b}\big[\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}j}^{b}\star\nabla^{(\ell)}\widehat{k}_{i}{}^{j}\big]+\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{c}\big[\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big]\\ -\frac{2}{3}\Phi g^{ab}g^{ii^{\prime}}g_{jj^{\prime}}\nabla_{i}\big[\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ab}^{j}\big]+g^{ii^{\prime}}g_{jj^{\prime}}g^{cj}\nabla_{i}\big[\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\big$$\bigg]$ (5.36)
Here the $\star$ symbol in all schematic expressions above signifies that all relevant indices in these terms are contracted.
@@
[…]
@@
[…]